OpenStudy (anonymous):

A salesman drives from Ajax to Barrington, a distance of 115 mi, at a steady speed. He then increases his speed by 14 mi/h to drive the 158 mi from Barrington to Collins. If the second leg of his trip took 8 min more time than the first leg, how fast was he driving between Ajax and Barrington?

OpenStudy (anonymous):

assuming steady rise in speed or instantly went from 14 m/h to x mi/h?

OpenStudy (anonymous):

i get 5.733 but I think I am wrong i'll dbl check

OpenStudy (anonymous):

now I get 46 for original speed :)

OpenStudy (anonymous):

sounds more reasonable 46 mi/h :)

OpenStudy (anonymous):

115=vt sfr t t=115/v here is where it gets complicated the second leg of trip is 158=(v+14)*(t + 2/15) 158=(v+14)*(115/v + 2/15) sfv =46

OpenStudy (anonymous):

@timo86m Thanks!!

OpenStudy (anonymous):

was it right?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

a cool I think that means he was going 46 + 14 the second leg of trip that is fast :P